(ii) The time for half life of first order reaction is 1 hr. What is time taken for  87.5% completion of the reaction. 
 

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
87.5 % means three half lives
As 87.5 % 
1st half life = 50%
2nd half life = 50 + 50% (50) = 50 +25 = 75%
3rd half life = 50 + 50 % (50) + 50% (25) = 87.5%
 
Hence total time taken = 3 hours
 
Regards
Arun (askIITians forum expert)